My question is regarding Sessa's reward. What is the equation to determine the total number of grains of rice after 64 days?
Answer by stanbon(75887)

My question is regarding Sessa's reward. What is the equation to determine the total number of grains of rice after 64 days?
------------
It's probably 1+2+2^2+2^3+2^4+2^5+ ...+ 2^63
---------------
S(n) = a[r^n-1]/[r-1] where r = 2
S(64)= 1[2^64 -1]/[2-1] = 2^64 -1 -1.844674407 x 10^19 grains
===============================================================
Cheers,
Stan H.